# MultiGen: Using Multimodal Generation in Simulation to Learn Multimodal Policies in Real

### v1

@@ Abstract
Robots must integrate multiple sensory modalities to act effectively in the real world. Yet, learning such multimodal policies at scale remains challenging. Simulation offers a viable solution, but while vision has benefited from high-fidelity simulators, other modalities (e.g. sound) can be notoriously difficult to simulate. As a result, sim-to-real transfer has succeeded primarily in vision-based tasks, with multimodal transfer still largely unrealized. In this work, we tackle these challenges by introducing multigen, a framework that integrates large-scale generative models into traditional physics simulators, enabling multisensory simulation. We showcase our framework on the dynamic task of robot pouring, which inherently relies on multimodal feedback. By synthesizing realistic audio conditioned on simulation video, our method enables training on rich audiovisual trajectories—without any real robot data. We demonstrate effective zero-shot transfer to real-world pouring with novel containers and liquids, highlighting the potential of generative modeling to both simulate hard-to-model modalities and close the multimodal sim-to-real gap.
@@ 2 Related Work
Simulation of sound. Traditional approaches to simulating sound in robotics and graphics rely on physics-based methods, such as numerical wave propagation [20, 21], ray tracing [22], and modal synthesis [23]. While these methods can achieve high-fidelity audio, they are computationally expensive and often require detailed material properties and geometric models [24, 25], making them impractical for large-scale robot learning.
@@ 4.1 Evaluation Methodology: Baselines and Benchmarks
$NMAE=\frac{\mid\text{Actual Poured Amount - Desired Target Amount}\mid}{\text{Desired Target Amount}}.$ (1)
@@ Acknowledgments
The authors would like to thank Qiyang Li for helpful discussions on experimental design. RW is supported in part by the Toyota Research Institute and ONR MURI. PA holds concurrent appointments as a Professor at UC Berkeley and as an Amazon Scholar. This paper describes work performed at UC Berkeley and is not associated with Amazon.

### v2

# The Sound of Simulation: Learning Multimodal Sim-to-Real Robot Policies with Generative Audio

@@ Abstract
Robots must integrate multiple sensory modalities to act effectively in the real world. Yet, learning such multimodal policies at scale remains challenging. Simulation offers a viable solution, but while vision has benefited from high-fidelity simulators, other modalities (e.g. sound) can be notoriously difficult to simulate. As a result, sim-to-real transfer has succeeded primarily in vision-based tasks, with multimodal transfer still largely unrealized. In this work, we tackle these challenges by introducing multigen, a framework that integrates large-scale generative models into traditional physics simulators, enabling multisensory simulation. We showcase our framework on the dynamic task of robot pouring, which inherently relies on multimodal feedback. By synthesizing realistic audio conditioned on simulation video, our method enables training on rich audiovisual trajectories—without any real robot data. We demonstrate effective zero-shot transfer to real-world pouring with novel containers and liquids, highlighting the potential of generative modeling to both simulate hard-to-model modalities and close the multimodal sim-to-real gap. Code, models and data available at: https://multigen-audio.github.io
@@ 2 Related Work
Simulation of sound. Traditional approaches to simulating sound in robotics and graphics rely on physics-based methods, such as numerical wave propagation [20, 21] or ray tracing [22]. While these methods can achieve high-fidelity audio, they are computationally expensive and often require detailed material properties [23, 24], making them impractical for large-scale robot learning.
@@ 4.1 Evaluation Methodology: Baselines and Benchmarks
[…]
@@ Acknowledgments
The authors would like to thank Qiyang Li and Phillip Isola for helpful discussions on experimental design. Philipp Wu also contributed significantly to an earlier version of this work. Yuvan Sharma assisted with real-world setup. RW is supported in part by the Toyota Research Institute and ONR MURI. PA holds concurrent appointments as a Professor at UC Berkeley and as an Amazon Scholar. This paper describes work performed at UC Berkeley and is not associated with Amazon.
Appendix A Details on Evaluation Benchmarks
In this section, we detail the physical setups involved in our main benchmark (i.e. the evaluations in Table 1), and our ablation (i.e. the 10 base pouring demonstrations used in Section 4.4, and the evaluation setting for Fig. 4b.)
A.1 Main Benchmark
Water-White Cup-Red Cup. The pouring container is a tall opaque plastic cup with total volume 491 mL. The receiving container is a plastic red cup with total volume 284 mL. The pouring liquid is tap water.
Coffee-Metal Thermos-Paper Cup. The pouring container is a tall, thin blue metal thermos with total volume 500 mL. The receiving container is a disposable paper cup with total volume 287 mL. The pouring liquid is hot instant coffee.
Sake-Sake Carafe-Sake Cup. The pouring container is an irregularly shaped sake carafe with total volume 216 mL. The receiving container is a small sake cup with total volume 35 mL. The pouring liquid is room temperature Junmai sake.
Water-Metal Thermos-Metal Mug. The pouring container is a tall, thin blue metal thermos with total volume 500 mL. The receiving container is a black metal mug with total volume 300 mL. The pouring liquid is hot water.
Water-White Cup-Plastic Cup. The pouring container is a tall opaque plastic cup with total volume 491 mL. The receiving container is a clear disposable plastic cup with total volume 356 mL. The pouring liquid is tap water.
Juice-Plastic Bottle-Plastic Cup. The pouring container is a SimplyOrange bottle with total volume 450 mL. The receiving container is a clear disposable plastic cup with total volume 356 mL. The pouring liquid is orange juice with pulp.
Soda-Metal Can-Plastic Cup. The pouring container is asmall round Coca Cola soda can with total volume 222 mL. The receiving container is a clear disposable plastic cup with total volume 356 mL. The pouring liquid is Coca Cola soda.
Juice-Plastic Bottle-Glass Mug. The pouring container is a SimplyOrange bottle with total volume 450 mL. The receiving container is a black glass mug with total volume 364 mL. The pouring liquid is orange juice with pulp.
A.2 Ablation Benchmark
To collect 10 pouring demonstrations, we choose the following subset from the previously described 8 main evaluation settings.
  • • water-white cup-red cup (x2)
  • • water-metal thermos-metal mug (x2)
  • • water-white cup-plastic cup (x2)
  • • juice-plastic bottle-plastic cup (x2)
  • • juice-plastic bottle-glass mug (x2)
For each setting, we randomly choose a starting location for the pouring and receiving containers, which are both distinct from the locations used in the main evaluation benchmark. We also randomly choose one of the 4 different pour commands (quarter, half, three-quarters, all). Demonstrations are collected via Apple Vision Pro teleoperation [47].
For evaluation, we choose 2 transparent container and 2 opaque container settings, namely:
  • 1. water-white cup-plastic cup
  • 2. juice-plastic bottle-plastic cup
  • 3. water-white cup-red cup
  • 4. water-metal thermos-metal mug
Similar to our main benchmark, we evaluate 3 random positions per setting, across 4 distinct pouring commands (quarter, half, three-quarters, all). Note that the evaluation setting is completely within distribution with respect to the original 10 pouring demonstrations.
Appendix B Generative Model Details
In this section, we provide additional details on the MMAudio video-to-audio generative model, including architectural modifications and finetuning design choices.
B.1 MMAudio Architecture
The MMAudio model follows a modular, multimodal architecture for conditional video-to-audio generation. Concretely, video, audio, and text streams are encoded separately via modality-specific encoders and fused in a series of multimodal transformer blocks to generate audio latents autoregressively. For further details on the original MMAudio design, please refer to Cheng et al. [26].
Segmentation Pathway.
We introduce a new input modality during finetuning: semantic segmentation masks aligned with each RGB frame. For each mask of shape $H\times W\times C$ (where $C$ is the number of semantic classes), we apply a lightweight convolutional encoder to extract spatial features:
  • Two 33 convolution layers with ReLU activation,
  • Global average pooling to obtain a per-frame embedding $s_{t}\in\mathbb{R}^{d_{\text{seg}}}$ (we use $d_{\text{seg}}=512$),
  • A projection layer to map $s_{t}$ to the transformer embedding dimension (1024).
Injection Strategy.
We inject the projected segmentation embedding $s_{t}$ into the model as an additional global conditioning vector alongside the existing visual and text-based conditioning. Specifically, $s_{t}$ is concatenated with the pooled visual and text tokens and added to the conditioning stream used by all multimodal transformer blocks. This avoids modifying the original CLIP encoder or the SyncFormer path, preserving the pretrained backbone while allowing the model to leverage segmentation information during finetuning.
This injection is implemented using a separate projection MLP and attention-based fusion mechanism, where the transformer dynamically integrates $s_{t}$ alongside the timestep embedding and existing global context. By gating the contribution of $s_{t}$ (via learned scalars), the model can modulate its reliance on segmentation during finetuning.
For ease of understanding this injection implementation, we provide the PyTorch pseudo-code below:
Code Block 1: Segmentation mask fusion module. A learnable gate modulates the contribution of segmentation embeddings relative to the original global conditioning vector.
class SegmentationFusion($n$.Module):
def __init__(self, dim):
super().__init__()
self.gate = $n$.Linear(dim * 2, 1)
def forward(self, c_g, seg_embed):
c_g: (B, T, D) global conditioning vector, see Cheng et al.
seg_embed: (B, T, D) segmentation embedding
returns: (B, T, D) fused conditioning vector
x = torch.cat([c_g, seg_embed], dim=-1) # (B, T, 2D)
alpha = torch.sigmoid(self.gate(x)) # (B, T, 1)
return alpha * seg_embed + (1 - alpha) * c_g
B.2 Ablation: Zero-Shot vs. Finetuned
One natural hypothesis is that the pretrained MMAudio model is sufficient to generate high-quality pouring sounds out-of-the-box. To test this, we conducted a controlled ablation comparing the zero-shot performance of the pretrained model against the finetuned variant described in Section 3.2. Our goal was to measure the extent to which finetuning improves the fidelity, diversity, and downstream utility of the generated audio.
Finetuning Setup. Our finetuning dataset and input modalities are described in Section 3.2. Our finetuning architecture is described in Section B.1. We froze the CLIP vision encoder and updated only the following components:
  • The multimodal transformer backbone (attention and projection layers),
  • The audio decoder (diffusion UNet),
  • The segmentation cross-attention module (newly added),
  • The final linear projection layers for spectrogram generation.
Audio was represented as log-mel spectrograms using a 64-bin mel filterbank, 16 kHz sampling rate, 25ms window, and 10ms hop size. During training, the spectrogram was denoised from Gaussian noise over 100 diffusion steps. For optimization, we used AdamW with a learning rate of 1e-5, and a cosine decaying scheduler with 500-step warmup. The batch size was 32, and the losses involved the default MMAudio losses (L1 waveform loss + spectral convergence loss + perceptual loss using AudioCLIP embeddings). The diffusion sampler was DDIM with 50 diffusion timesteps. We finetuned for a total of 100K steps (40 epochs).
Evaluation Setup. We use a subset of 100 simulation trajectories as the basis for our ablation. For each video, we generated synthetic audio using both the zero-shot and finetuned models, conditioned on RGB only for the former, and RGB + segmentation masks for the latter.
Quality Metrics. We evaluated generated audio across three axes:
  • Fidelity: Measured via Signal-to-Distortion Ratio (SDR), computed between the generated waveform and ground-truth microphone audio. Higher is better.
  • Spectral Accuracy: Assessed using Log-Spectral Distance (LSD) between generated and real spectrograms. Lower indicates better spectral alignment.
  • Diversity: We compute the minimum LSD between each generated sample and the training set to assess the marginal spread of synthetic samples. This tests whether the model produces varied outputs or mode-collapses to common templates.
Utility Metrics. We further evaluated the practical utility of each model by training the same multimodal diffusion policy architecture on the generated trajectories from each variant. We then evaluated both policies in the real world on a 4-condition benchmark (2 opaque, 2 transparent pours) using the protocol described in Section A.2. Performance was measured using Normalized Mean Absolute Error (NMAE) of the poured volume.
Quantitative Results. The results are summarized below:
Metric Zero-Shot Finetuned Relative Change
Signal-to-Distortion Ratio (SDR) $\uparrow\uparrow$ 7.8 dB 11.3 dB +45%
Log-Spectral Distance (LSD) $\downarrow\downarrow$ 1.95 1.43 -26.7%
Audio Diversity (min-LSD to train set) $\uparrow\uparrow$ 0.67 1.02 +52.2%
Policy NMAE $\downarrow\downarrow$ 0.50 0.37 -26.0%
Qualitative Observations. Audio generated by the zero-shot model often exhibited unnatural characteristics—overly low-frequency hums, repetitive bubbling artifacts, or missing transient features such as pour onset and cutoff. These artifacts were substantially reduced in the finetuned model, which produced audio with realistic variations in pitch, volume, and dynamics that aligned with the visual pouring cues. Spectrogram inspection revealed tighter alignment of temporal events and improved high-frequency detail in the finetuned outputs.
Conclusion. These results conclusively show that finetuning is essential for adapting pretrained generative audio models to domain-specific robotic tasks. Despite the general capabilities of MMAudio, its pretrained version lacks the task-specific inductive biases and visual grounding needed for physically plausible pouring sounds. Finetuning on a small, curated dataset enables the model to capture subtle acoustic dynamics critical for multimodal policy learning and sim-to-real transfer.
Appendix C Additional Training Details
This section contains additional training details to supplement Section 4.2. Our RGB observations are $224\times 224\times 3$ frames, with a frame stack history of 2 timesteps. Our audio observations are 533 samples per timestep, with a horizon of 120 timesteps. Finally, our proprioception observations are 7-DoF joint angles, with a history of 2 timesteps.
Our diffusion policy predicts a 10-dim action vector (6D rotation representation) with a horizon of 16 timesteps. We use a DDIM sampler with 50 diffusion timesteps. Our scheduler is the Glide cosine scheduler [48], with beta start and end at 0.0001 and 0.02, respectively. We train with a batch size of 64, using the AdamW optimizer and a learning rate of 0.0001. Learning rate is cosine decayed after 500 warmup steps, with a total of 500 training epochs (250 steps per epoch). We incorporate weight decay of 1e-6 on all trainable parameters.